OpenStudy (anonymous):

The following lines are parallel. 2x + 2y = 4 and y = -x + 5 True or false?

OpenStudy (mertsj):

Do they have the same slope?

OpenStudy (lgbasallote):

turn 2x + 2y = 4 into slope intercept form 2y = -2x + 4 now divide everything by 2

OpenStudy (anonymous):

y=x+3

OpenStudy (lgbasallote):

no...it's y = -x + 2 if yu divide everything by 2

OpenStudy (anonymous):

oh haha messed up there.

OpenStudy (lgbasallote):

now look at the coefficient of x of both equation...are they the same?

OpenStudy (anonymous):

nope?

OpenStudy (anonymous):

wait yes

OpenStudy (lgbasallote):

the x are the same therefore they are parallel

OpenStudy (anonymous):

Thanks! Now I know how to find them the next time! :)